# Population of tetraneutron continuum in reactions of 8He on deuterium

### v5

@@ Abstract
Search for the population of the low-energy continuum of tetraneutron system was performed for the 2H(8He,6Li) and 2H(8He,3He)7HH+ reactions. Evidence for a hump in the continuum at and MeV, respectively, was observed in both reactions. The observed statistics is quite low (6 and up to 40 events) corresponding to very low cross sections of few microbarns or tens of microbarns. The background conditions for the 2H(8He,6Li) reaction are shown to be good, favoring the physical nature of the observed events. The 2H(8He,3He)7HH+ process transforms to the 2H(8He,6Li reaction in the limit of the highest 7H decay energies. The population of the low-energy region in the 4n spectrum is found to be perfectly correlated with the population of the lowest 6Li states in the 3He+3H continuum. Theoretical calculations of 8He in a five-body + and of 4n in a four-body hyperspherical models are presented. The 8He wave function is shown to contain strong specific correlations, which may give rise to very low-energy structures in tetraneutron continuum in extreme-peripheral reaction scenarios.
Keywords: tetraneutron spectrum; 8He beam; direct reactions; cryogenic deuterium target; hyperspherical harmonics; five-body + model for 8He; four-body continuum.
Date: April 30, 2024.
1 Introduction
@@
The search for the multineutron systems is old, but still unsettled problem of the low-energy nuclear physics. The first ideas about the possible multineutron systems stability were expressed in the papers [1, 2] and since that time reiterated on various occasions. A detailed account of the multineutron studies history, both experimental and theoretical was provided in the recent review [3]. In short, either none or only marginal experimental evidence, which was never confirmed later, was obtained for the bound tetraneutron. However, such attempts continue, and there is a very recent example [4]. The situation has changed with the recent studies of 4n population in reactions with 8He, where four neutrons can be found in a spatially-separated neutron-halo configuration. In the 4He(8He,) reaction [5] 4 events were observed in the low center-of-mass (c.m.) 4n energy window MeV with the tiny cross section of nb. Even more recently, a statistically convincing peak (reported as “resonance-like structure”, events) was observed in a 1H(8He,) experiment [6] at MeV with MeV. The reported cross section in this study is also small and estimated to be sub-microbarn.
@@
The issue of bound “multineutron nuclei” was scrutinized in the modern theoretical approaches [7, 8, 9] always with the same result: with fixed two-body potentials, radical modifications of the three-body potential are required. Such modifications are considered unacceptable as they are inconsistent with well-known nuclear structure. In the theoretical studies of the 4n low-energy continuum various indications were found for important low-energy effect of the 4n final-state interaction (FSI). These include: specific trajectories of the -matrix poles, enhanced “time delay”, energy extrapolations of multi-neutron states confined in traps, etc., see e.g. [10, 11, 12, 13, 14, 15, 16] and, especially, the illuminating summary of those studies in [3]. There have been two works that, unlike those cited above, have predicted a 4n resonant state [17, 18], strongly contradicting the other studies.
@@
An approach to 4n system tractable in terms of observables (namely, the reactions with neutron-rich halos) was presented in [19], wherein the 4n continuum is populated from the “prearranged” 4 neutron configurations in the “atmosphere” of a halo nucleus like 8He. The approach predicts population at low energies, but this population is a result of a specific (very peripheral) initial state structure (ISS) sampling, modified by FSI and should not be considered a stand-alone resonance. The formation of such a low-energy continuum response is partly analogous to the mechanism of the “soft dipole mode” formation in halo nuclei: (i) relatively very weak FSI effects and (ii) strong low-energy concentration of the strength function due to the peripheral character of the halo wave functions (WF) and the periphery enhancement by electromagnetic transition operators [20, 21, 22]. Principal message of [19] is that both halo character of 8He WF and 4n FSI are not sufficient to produce low-energy responce in the 4n continuum, additional peripheral assumption about reaction mechanism is required. In [23] the data of [6] were analyzed under assumptions similar to Ref. [19]. However, conclusion was made in [23] that the 8He-induced source can directly explain the data of [6] without any peripheral assumptions. This is in a qualitative contradiction with the results of [19].
@@
In this work we demonstrate that an evidence for the low-energy structures analogous to the observation of [5, 6] can be found in the other reactions with the 8He beam. As for the theoretical discussion, we revisit the 4n population from 8He in simple approach [19], but important technical improvements. We generally confirm the results of [19], but provide an additional insight into the reaction mechanism, connected with specific correlations in the 8He WF. Possible reasons for disagreement with the results of [23] are analyzed. The system of units is used in this work.
2 Experiment
@@
Experiment with the 8He beam impinging the deuterium target was performed at the ACCULINNA-2 facility (FLNR, JINR), having the search for the 7H nuclide as a goal [24, 25]. The experiment appeared to be much “richer” than intended, providing interesting results also on 6H [26] and other auxiliary channels like 8Li, 9Li [27, 28] and 5H, 5He, 7He [29]. Here we present the results obtained for the 2H(8He,6Li) and 2H(8He,3He)7HH+ reactions, previously omitted. Since the experiment has been already well presented in literature, we only briefly describe here the most relevant details.
@@
The 8He beam was produced at 26 A MeV with intensity pps. The beam was focused on the cryogenic gaseous deuterium target, with a temperature of 27 K, equipped with the thin stainless-steel and mylar windows. The target thickness was . The setup allowed us to measure the 2H(8He,6Li)4n reaction and some products of the multibody decay of the produced unbound system.
@@
The beam particles were identified by the two plastic scintillators, allowing to deduce the energy of the projectile from its measured time-of-flight (ToF). The projectile trajectories were reconstructed by the two pairs of multi-wire proportional chambers. The special run with the empty target cell was performed to estimate the background conditions, which had of the total 8He beam time.
@@
The recoil 6Li nuclei, appearing in the 2H(8He,6Li) reaction hit the array of four identical -- telescopes. The telescope array was located 179 mm downstream the target. Each telescope consisted of three layers of silicon strip detectors (SSDs). The 20-m-thick SSD with a sensitive area of mm2 was divided into 16 strips, the second and the third layers were created by the two identical 1 mm-thick SSDs ( mm2 with 16 strips). The 6Li recoils emitted from the deuterium gas target in the 2H(8He,6Li) reaction in a range degrees in the lab system were detected by this telescope array. The telescopes allowed to identify 6Li with clear separation from the other registered lithium isotopes, see Fig. 1 (a). The central telescope was installed at the beam line at the distance of 323 mm behind the target. It was intended to detect tritons emitted with high energies at the angles smaller than in the laboratory system. The telescope consisted of one 1.5-mm-thick double-sided SSD ( mm2, with 32 strips on each side) followed by a square array of 16 CsI(Tl) crystals. The crystals had a cross section of mm2 and thickness 50 mm each.
@@
The group of four neutrons appears in “free flight” as a result of the -core removal from the 8He projectile. An important part of the 2H(8He,6Li)4n reaction analysis was the neutron identification and reconstruction. The neutron-wall setup [30] included 48 stilbene scintillator crystals placed on a m2 area located 2 m downstream the target at zero angle to the beam axis. The distance between the 50-mm thick and 80-mm diameter stilbene crystals was approximately 12 cm which resulted in about probability for neutrons to hit a stilbene detector. The stilbene array provided energy resolution and the single neutron registration efficiency of . The probability of a neutron registration in coincidence with the 6Li recoil was around , taking into account that four neutrons are flying forward, towards the stilbene array. The -gamma separation was made by means of the Pulse Shape Discrimination (PSD) [30]. The PSD information, supplemented with ToF distribution, see Fig. 1 (b), leads to suggestion that some gamma-type signals correspond to neutron-like ToF. These gammas are presumably produced by neutrons interacting with the stilbene modules housings and should be taken as neutron events.
2.1 2H(8He,6Li)4n results
@@
Tetraneutron was reconstructed from the recoil 6Li as a missing-mass (MM) component in the 2H(8He,6Li) reaction. The total number of 6Li- coincidences found in the recorded data was 136, see Fig. 2 (a). In this figure the neutron kinetic energy in the 4n c.m. frame is compared with the reconstructed MM energy of the group. The fact that the majority of these events (108 events) are located inside the kinematically allowed region proves a good channel identification for the data. The empty-target measurement Fig. 2 (a) shows that the background conditions were very “clean” for the 4n population in the 2H(8He,6Li) reaction: the “empty” events are very few and located mainly in the unphysical part of the kinematical plane.
@@
The MM spectrum of 4n, Fig. 2 (b), was reconstructed from the measured 6Li recoil energy and emission angle, taking the 108 events located within the “kinematical triangle” . The obtained spectrum shows a group of 6 events around MeV. The width of this group is well consistent with the value of energy resolution MeV FWHM in this region of 4n MM spectrum. The statistics of this low-energy group is qite low, but the energy resolution is sufficient and the background conditions are shown above to be very good. The green dotted curve Fig. 2 (b) shows that less that 1 random event is expected under MeV. Then the Poisson probability to get 6 random coincidences in that energy region is then .
Obviously, the observed low-energy hump in Fig. 2 (b) can not be described just as a contribution of the 4-body phase volume or “Fourier transform” of the 8He source, see also Fig. 8 (d). The “generic” 4-body phase volume behavior is , but for the 4n population the Pauli principle requires at least two additional “excitation quanta” for the lowest-energy configuration or in terms of the hyperspherical harmonics (HH) model, see Section 4. This leads to a modified 4-body phase-volume behavior . It is notable that the observed energy of this group events (sys.) MeV is consistent within the experimental energy resolution with the 4n low-energy peak value (sys.) MeV reported in Ref. [6].
2.2 2H(8He,3He)7HH+ results
@@
In Ref. [25] the excitation spectrum of 7H was populated in the 2H(8He,3He)7H reaction up to MeV above the 5-body 3H+ decay threshold. An evidence was reported for resonant states at and 5.5(3) MeV and also some indications for such states at 7.5(3) and 11.0(3) MeV. These structures were observed at small c.m. reaction angles . The 7H excitations in a broad range with MeV form quite a featureless hump up to maximal available energy. It is possible to search for the low-energy-correlated 4n emission off these sufficiently highly-excited configurations of 7H.
@@
The data of Ref. [25] contain coincidence information on the 3H decay product of 7H, which allows one to infer the MM of the 4n subsystem. These data in terms of the 4n emission are shown in Fig. 3 (a). The 7H events are mainly concentrated close to diagonal , which means that the “internal” energy of the tetraneutron tends to be large, while the triton gets only a small fraction of the total 7H decay energy. This situation is close to the phase-volume distribution of the 5-body 3H+ decay for the given 7H decay energy as
@@
in terms of 4n energy-distribution parameter . The standard value for the 5-body phase volume is . It was discussed in the previous section that at least value is requested for by the Pauli principle for 4n. This idea is confirmed by the calculations of the 5-body decay of 7H in [31]. The distributions at given 7H energies are also quite “compressed” due to the experimental bias, see Fig. 13 of [25], so for the distribution of Eq. (1) turns out to be completely justified in our situation.
@@
To observe a possible 4n correlation, we need to find the kinematical region, where the main “phase-volume” component of the decay distribution Eq. (1) is well suppressed. This means , and for the studies of the 4n excitation region MeV we need to consider the 7H decay energy region MeV. Under this selection condition, the hump with MeV can be found in the 4n MM spectrum, see Fig. 3 (b). The Monte Carlo (MC) simulated phase volume only weakly “penetrates” in the MeV region. The background of about events/MeV is estimated from the available empty target data, see green circles in Fig. 3 (b). Thus a significant part of this low-energy 4n data may be connected (but not necessarily is connected) with the random background.
@@
The 7H data can be converted to presentation of system reconstructed as 3He+3H continuum, see Fig. 4 (a). Fortunately, it appeared that the low-energy 4n events in Fig. 3 (b) are practically perfectly correlated with the population of the lowest state in the 3He+3H continuum, with MeV and MeV [32], see Fig. 4 (b). This means that we are actually dealing here not with the 7HH+ decay, but rather with the 2H(8He,6Li reaction. Moreover, the same correlation is perfectly true not only for events in the MeV range, but also for the MeV range. In this range most of the events should be either unphysical or connected with insufficient energy resolution, as the population of the MeV range is predicted to be negligible in available theoretical scenarios, see Fig. 7 or Ref. [23]. The MC estimated energy resolution for the 4n MM spectrum is MeV FWHM. However, this resolution has a broad “nongaussian” component, which may result in some number of events with MeV. The coincidence with the definite state in 6Li practically guarantee the physical nature of such event, so we recognize that the absolute majority of events (26 out of 28) found in the range MeV in Fig. 3 (b) as really belonging to the 4n MM spectrum.
@@
Strong correlation found in Fig. 4 (b) inspired us to “reverse” the logic and construct the 4n spectrum gated by the lowest states in the continuum, see Fig. 4 (c). The profile of the low-energy MeV 4n spectrum is practically insensitive to the gate energy selection for MeV. For higher energies the 4n spectrum expectedly becomes phase-volume-like. The peak energy can be found as (sys.) MeV. That peak position is consistent with (sys.) MeV of Ref. [6] within the experimental energy resolution. Statistically our data is analogous to data [6]: 44 events in [6] vs. 26 and 40 events in the and MeV gates, respectively.
@@
There is a strong sensitivity of the obtained 4n MM spectrum to the energy of the populated 6Li states. For example, it can be seen in Fig. 4 (c) that 4n spectra in the gates and MeV show populations 4 events versus 25 events in the MeV range, which seem to be beyond statistical uncertainty. This may be indication that mechanisms for 4n population within these gates are somewhat different: the states of the 6Li recoil are broad and overlappind in this energy range [32], and the higher-energy gate could be “contaminated” by contributions of the higher excitations of 6Li.
3 On 4n population in different reactions

### v8

@@ Abstract
Search for the population of the low-energy continuum of a tetraneutron system was performed for reactions of the 8He beam on a deuterium target. These studies are based on the data [I.A. Muzalevskii et al., Phys. Rev. C 103, 044313 (2021)], previously used for the studies of 7H and 6H in the H and H reactions. Evidence for a hump in the 4n continuum at and MeV was observed in the 2H(8He,6Li)4n and 2H(8He,3He)7HH+4n reactions, respectively. The observed statistics is quite low (6 events and up to 40 events) corresponding to very low cross sections of few microbarns or tens of microbarns. The background conditions for the 2H(8He,6Li)4n reaction are shown to be good, favoring the physical nature of the observed events. The 2H(8He,3He)7HH+4n process transforms to the 2H(8He,6Li reaction in the limit of the highest 7H decay energies. The population of the low-energy region in the 4n spectrum is found to be perfectly correlated with the population of the lowest 6Li state in the 3He+3H continuum with MeV. Theoretical calculations of 8He in a five-body + and of 4n in a four-body hyperspherical models are presented. The 8He wave function is shown to contain strong specific correlations, which may give rise to very low-energy structures in tetraneutron continuum in extreme-peripheral reaction scenarios.
I Introduction
@@
The search for multineutron systems is old, but still unsettled problem of the low-energy nuclear physics. The first ideas about the possible nuclear-stability of multineutron systems were expressed in Refs. [1, 2] and since that time reexamined on various occasions. A detailed account of the search for bound multineutron, both experimental and theoretical, can be found in the recent review [3]. In short, either none or only marginal experimental evidence, which was never confirmed later, was obtained for the bound tetraneutron. However, such attempts continue, and there is a very recent example [4]. The situation has changed with the recent studies of the low-energy 4n continuum population in reactions with 8He, where four neutrons can be found in a spatially-separated neutron-halo configuration. In the 4He(8He,) reaction [5] 4 events were observed in the low center-of-mass (c.m.) 4n energy range MeV with the tiny cross section of nb. Even more recently, a statistically convincing peak (reported as “resonance-like structure” comprising around 44 events) was observed in a 1H(8He,) experiment [6] at MeV with MeV. The reported cross section in this study is also small and estimated to be sub-microbarn. It should be noted that the both observations [5, 6] belong to the very specific situation of extreme backward (backward/forward in the case of [5]) kinematics of quasi-free scattering.
@@
The issue of bound “multineutron nuclei” was scrutinized in the modern theoretical approaches [7, 8, 9] with the same result: for fixed two-body potentials, radical modifications of the three-body potential are required to get bound 4n. Such modifications are considered unacceptable as they are inconsistent with well-known nuclear structure of all other nuclides. In the theoretical studies of the low-energy 4n continuum various indications were found for important low-energy effect of the 4n final-state interaction (FSI). These include: specific trajectories of the -matrix poles, enhanced “time delay”, energy extrapolations of multi-neutron states confined in traps, etc., see e.g. [10, 11, 12, 13, 14, 15, 16]. There have been, however, two works that have predicted a 4n resonant state [17, 18], strongly contradicting the other studies. It has been mentioned above that the illuminating summary of those studies can be found in the review [3].
@@
An approach to 4n system tractable in terms of observables (namely, the reactions with neutron-rich halos) was presented in Ref. [19]. In this work the 4n continuum is populated from the “prearranged” 4 neutron configurations in the “atmosphere” of a halo nucleus 8He. The approach predicts population at low energies, but this population is a result of a specific (very peripheral) initial state structure (ISS) sampling, modified by FSI and should not be considered a stand-alone resonance. The formation of such a low-energy continuum response is partly analogous to the mechanism of the “soft dipole mode” formation in halo nuclei: (i) relatively very weak FSI effects, but nevertheless, strong low-energy concentration of the strength function due to (ii) the peripheral character of the halo wave functions (WF) and (iii) the further periphery enhancement by electromagnetic transition operators [20, 21, 22]. Principal message of [19] is that both halo character of 8He WF and 4n FSI are not yet sufficient to produce low-energy response in the 4n continuum, and additional assumption about peripheral nature of the reaction mechanism is required.
In Ref. [23] the data of [6] were analyzed under assumptions similar to Ref. [19]. However, conclusion was made in [23] that the 8He-induced source can directly explain the data of [6] without any peripheral assumptions on reaction mechanism. So, there is a qualitative contradiction between the results of [19] and [23] based on more or less the same physical assumptions.
@@
In this work we demonstrate that an evidence for the low-energy structures, analogous to the observations of [5, 6], can be found in the other reactions with the 8He beam. As for the theoretical discussion, we revisit the 4n population from 8He in simple reaction approach of Ref. [19] (sudden removal approximation) introducing, however, important technical improvements. We generally confirm the results of [19], but provide an additional insight into the reaction mechanism, connected with specific correlations in the 8He WF. Possible reasons for disagreement with the results of [23] are analyzed.
The system of units is used in this work.
II Experiment
@@
Experiments with the 8He beam impinging on the deuterium target was performed at the ACCULINNA-2 facility (FLNR, JINR) having the search for the 7H nuclide as a goal [24, 25]. The experiment appeared to be much “richer” than initially intended, providing interesting results also on 6H [26] and other auxiliary channels like 8Li, 9Li [27, 28] and 5H, 5He, 7He [29]. Here we present the results obtained for the 2H(8He,6Li) and 2H(8He,3He)7HH+4n reactions, previously omitted. Since the experiment has been already well presented in literature, we only briefly describe here the most relevant details.
@@
The 8He beam was produced at 26 A MeV with mean intensity around pps. The beam was focused on the cryogenic gaseous deuterium target, with a temperature of 27 K, equipped with the thin stainless-steel and mylar windows. The target thickness was . For the 2H(8He,6Li)4n reaction the setup allowed us to measure the 6Li recoil in coincidence with a neutron from the produced unbound 4n system.
@@
The beam particles were identified by the two plastic scintillators, allowing to deduce the energy of the projectile from its measured time-of-flight (ToF). The projectile trajectories were reconstructed by the two pairs of multi-wire proportional chambers. A special run with empty target cell was performed to estimate the background conditions, which was of the total 8He beam time long.
@@
The recoil 6Li nuclei, appearing in the 2H(8He,6Li) reaction hit the array of four identical -- telescopes. The telescope array was located 179 mm downstream the target. Each telescope consisted of three layers of silicon strip detectors (SSDs). The 20-m-thick SSD with a sensitive area of mm2 was divided into 16 strips, the second and the third layers consisted of the two identical 1 mm-thick SSDs ( mm2 with 16 strips). The 6Li recoils emitted from the deuterium gas target in the 2H(8He,6Li) reaction were detected by this telescope array in a range in the lab system. The telescopes allowed to identify 6Li with clear separation from the other registered lithium isotopes, see Fig. 1 (a). The central telescope was installed at the beam line at the distance of 323 mm behind the target. It was intended to detect tritons emitted with high energies at the angles smaller than in the lab system. The telescope consisted of one 1.5-mm-thick double-sided SSD ( mm2, with 32 strips on each side) followed by a square array of 16 CsI(Tl) crystals. The crystals had a cross section of mm2 and thickness 50 mm each.
@@
The group of four neutrons appears in “free flight” as a result of the -core removal from the 8He projectile. An important part of the 2H(8He,6Li)4n reaction analysis was the neutron identification and reconstruction. The neutron-wall setup [30] included 48 stilbene scintillator crystals placed on a m2 area located 2 m downstream the target at zero angle to the beam axis. The distance between the 50-mm thick and 80-mm diameter stilbene crystals was approximately 12 cm which resulted in about probability for neutrons to hit a stilbene detector. The stilbene array provided energy resolution and the cumulative single-neutron registration efficiency of . The probability of a neutron registration in coincidence with the 6Li recoil can be estimated as , taking into account that four neutrons are flying forward, towards the stilbene array. The -gamma separation was made by means of the Pulse Shape Discrimination (PSD) [30]. The PSD information, supplemented with ToF distribution, see Fig. 1 (b), leads to suggestion that some gamma-type signals correspond to neutron-like ToF. These gammas are presumably produced by neutrons interacting with the stilbene modules housings and should be taken as neutron events.
II.1 2H(8He,6Li)4n results
@@
Tetraneutron spectrum was reconstructed from the recoil 6Li as a missing-mass (MM) component in the 2H(8He,6Li) reaction. For background reduction we considered 6Li coincidences with one of the neutrons coming from the 4n decay. The total number of 6Li- coincidences found in the recorded data was 136, see Fig. 2 (a). In this figure the neutron kinetic energy in the 4n c.m. frame is plotted opposite the reconstructed MM energy of the group. The fact that the majority of these events (108 events) are located inside the kinematically allowed region confirms a good channel identification for the data. The empty-target measurement Fig. 2 (a) shows that the background conditions were very “clean” for the 4n population in the 2H(8He,6Li) reaction: the empty-target events are very few and located mainly in the unphysical part of the kinematical plane.
@@
The MM spectrum of 4n, Fig. 2 (b), was reconstructed from the measured 6Li recoil energy and emission angle, taking the 108 events located within the “kinematical triangle” . The obtained spectrum shows a group of 6 events around MeV. The width of this group is reasonably consistent with the value of energy resolution MeV FWHM in this region of 4n MM spectrum. The green dotted curve in Fig. 2 (b) shows the Monte Carlo (MC) simulated empty-target contribution. The statistics of the low-energy group of events is quite low, but the energy resolution is sufficient and the background conditions are shown in Fig. 2 to be very good. It can be evaluated that less than 1 random event is expected in the energy range MeV. Then the Poisson probability to get 6 random coincidences in that energy region is .
Obviously, the observed low-energy hump in Fig. 2 (b) can not be described just as a contribution of the 4-body phase volume or Fourier transform of the 8He source, see also Fig. 8 (d) and the related discussion in Sec. VI. The phase volume for uncorrelated decay of -body system, which can be regarded here as a “minimum admissible model”, is
and the generic 4-body phase volume behavior is . However, Eq. (1) should be modified by the minimum number of excitation quanta allowed for the continuum WF as
For the 4n system the Pauli principle requires two additional excitation quanta for the lowest-energy configuration. This corresponds to lowest-energy configuration in the shell-model terms or to in terms of the hyperspherical harmonics (HH) model principal quantum number (hypermomentum), see Section IV. This leads to a modified 4-body phase-volume behavior shown in Fig. 2 (b).
It is notable that the mean energy MeV of the low-energy group of events in Fig. 2 (b) is consistent (within experimental energy resolution) with the 4n low-energy peak value MeV reported in Ref. [6].
The 2H(8He,6Li)4n reaction has already been used before for tetraneutron search, but the results were not published in details, only briefly in conference proceedings [31, 32]. In Ref. [31], a resonance-like structure (with quite a low statistics) was reported at MeV. In contrast, for the next analogous experimental run [32] the 4n spectrum with MeV was found to be dominated by the carbon background due to the use of the CD2 target. As compared to studies of [31, 32] our experiment has no carbon background and additional strong background elimination, connected with kinematics of Fig. 2 (a), is possible.
Another background issue to be discussed is possibility that the 2H(8He,6Li)4n reaction results could be contaminated by the 2H(8He,n reaction contribution, where is the isobaric-analogue state of the 6He(g.s.). Since the state decays only electromagnetically, we cannot distinguish the latter reaction contribution in the current experiment.
Fortunately, there is a strong argument why the population of should be very small. It is well established that reactions at typical energies of our experiment (26 ) and at higher energies are dominated by single-step direct-reaction mechanism. Within this paradigm the 4n subsystem is “spectator” (inert during the reaction time) and transfer is induced by the T-matrix between the 4He “participant” and the deuterium of the target:
where is sum of all possible nuclear interactions in the channel of interest. However, it is well known that the state can be obtained in 2H capture on -particle only by electromagnetic M1 transition or by parity-non-conservation neutral-current weak interaction:
where the parity-violating nuclear interaction is known to be extremely small. Indeed, the cross section for the capture reaction 4He(,Li has been measured in a broad energy range by detecting gammas or recoil 6Li ions [33]. No 3.56 MeV resonance population has been observed and the best limit on the parity-forbidden width of this state was obtained as tiny eV in Ref. [34]. Thus, the 2H(8He,n reaction may be connected only with two-step and more complicated reaction scenarios, which are expected to be suppressed.
It should be also understood that actual population will not dismiss the low-energy 4n issue, just make its interpretation more complicated. Assume extreme situation that only was actually populated. This means that we miscalculate the missing mass of 4n, which should be actually 3.56 MeV smaller according to energy-conservation. So, the low-energy 4n events in Fig. 2, which have energies between 2.5 and 5 MeV after recalculation will have energies from to 1.5 MeV with centroid at around the 4n threshold. Existence of such a “severe low-energy” 4n correlation, is even less realistic and complicated to interpret theoretically.
II.2 2H(8He,3He)7HH+ results
@@
In Ref. [25] the excitation spectrum of 7H was populated in the 2H(8He,3He)7H reaction up to MeV above the 5-body 3H+ decay threshold. An evidence was reported for resonant states at and 5.5(3) MeV and also some indications for such states at 7.5(3) and 11.0(3) MeV. These structures were observed at small c.m. reaction angles . The 7H excitations with MeV observed in a broad angular range form quite a featureless hump. It is possible to search for the low-energy-correlated 4n emission off these sufficiently highly-excited configurations of 7H.
@@
The data of Ref. [25] contain coincidence information on the 3H decay product of 7H, which allows one to infer the MM of the 4n subsystem. These data in terms of the 4n emission are shown in Fig. 3 (a). The 7H events are mainly concentrated near the diagonal , which means that the “internal” energy of the tetraneutron tends to be large, while the triton gets only a small fraction of the total 7H decay energy. This situation is close to the phase-volume distribution of the 5-body 3H+ decay for the given 7H decay energy as
in terms of 4n energy-distribution parameter
@@
The standard value for the 5-body phase volume is . Eq. (3) is the double-differential form of the Eq. (2), discussed in the previous Section, so it can be found that at least value is requested for by the Pauli principle for 4n. This idea is confirmed by the calculations of the 5-body decay of 7H in [35]. The distributions at given 7H energies are also quite “compressed” towards small 3H energies due to the experimental bias, see Fig. 13 of [25], so for the distribution of Eq. (3) turns out to be completely justified in our situation.
@@
To observe a possible low-energy 4n correlation, we need to find kinematic region, where the main “phase-volume” component of the decay distribution Eq. (3) is well suppressed. This means , and, therefore, for studies of the 4n excitation region MeV we need to consider the 7H decay energy region MeV. Under this selection condition, the hump with MeV can be found in the 4n MM spectrum, see Fig. 3 (b). The MC simulated phase volume only weakly “penetrates” in the MeV region. The background of about events/MeV is estimated from the available empty target data, see green circles in Fig. 3 (b). Thus a significant part of this low-energy 4n data may be connected (but not necessarily is connected, as we see below) with a random background.
@@
The 7H data can be converted to presentation of system reconstructed as 3He+3H continuum, see Fig. 4 (a). Fortunately, it appeared that the low-energy 4n events in Fig. 3 (b) are practically perfectly correlated with the population of the lowest state in the 3He+3H continuum, with MeV and MeV [36], see Fig. 4 (b). This means that we are actually dealing here not with the decay, but rather with the 2H(8He,6Lin reaction. Moreover, the same correlation is perfectly true not only for events in the MeV range, but also for the MeV range. In this range most of the events should be either unphysical or connected with insufficient energy resolution, as the population of the MeV range is predicted to be negligible in available theoretical scenarios, see Fig. 7 or Ref. [23]. The MC estimated energy resolution for the 4n MM spectrum is around MeV FWHM. However, this resolution has a broad “nongaussian” component, which may result in some number of events with MeV. The coincidence with the definite state in 6Li practically guarantee the physical nature of such event, so we recognize that the absolute majority of events (26 out of 28) found in the range MeV in Fig. 3 (b) as really belonging to the 4n MM spectrum.
@@
Strong correlation found in Fig. 4 (b) inspired us to “reverse” the logic and construct the 4n spectrum gated by the lowest states in the He+3H continuum, see Fig. 4 (c). The profile of the low-energy MeV 4n spectrum is practically insensitive to the gate energy selection for MeV. For higher energies the 4n spectrum expectedly becomes phase-volume-like. The peak energy can be found as (sys.) MeV. That peak position is consistent with (sys.) MeV of Ref. [6] within the experimental energy resolution. Statistically our data is analogous to data [6]: 44 events in [6] vs. 26 and 40 events in the and MeV gates, respectively.
@@
There is a strong sensitivity of the higher-energy part of the obtained 4n MM spectra to the energy of the populated 6Li states. For example, it can be seen in Fig. 4 (c) that 4n spectra in the gates and MeV show very different populations of 4 and 25 events in the MeV range, which seems to be beyond statistical uncertainty. This may be indication that mechanisms for 4n population within these gates are somewhat different: the states of the 6Li recoil are broad and overlapping in this energy range [36], and the higher-energy gate could be “contaminated” by contributions of the higher excitations of 6Li.
III On tetraneutron population in different reactions